The density of dry air at 0.9866 bar and 27°C is 1.146 g/dm³. Calculate the percentage nitrogen by mole.
1
Expert's answer
2020-11-22T11:11:52-0500

PV = nRT

0.9866 bar to atm = 0.974atm

27°C = 300K

n= PV/RT = 0.974 ×1 / (0.0821×300) = 0.040mole


Percentage by mole = 78/100 × 0.040 = 0.0312